3 mole of A and 2 mole of B are mixed with 4 litre vessel reaction 3A+2B----4C at equilibrium the flask contain 1 moles of B. What is kC

Grade:12th pass

4 Answers

Arun
4 years ago
Dear student
 
3A             +         2B.        =            4C
(3 – 3x ).           2 – 2x.                     4x
 
I think now you can solve.
 
In case of any difficulty,plfeel free to ask gain
yohib ahmed
4 years ago
Please answer in details if you can... please because i already 2-3 steps but answer not match, answer of this question is 18.97M-1
 
Arun
4 years ago
2 – 2x = 1
 
hence x = ½
 
Kc = (2/4)^4 / (3/8)^3 (1/4)^2 = 18.963
 
now Kc = 18.963
 
Hope it helps
 
Regards
Arun
yohib ahmed
4 years ago
Thanks Done Bro. your first step is very helpful for me. Thank you.
 
 
 Answer this question also please
 
Equilibrium constant of the reaction A+B ----- C+D is 1/3. How many moles of "A" must be mixed with 3 moles of "B" to yield at the equilibrium 2 mole of "C". The volume of the vessel is 2 litres 

Please solution this question.. 
Answer is 14 moles